### v1

# Energy Extraction via Magnetic Reconnection in Kerr-Sen-AdS4 Black Hole: Circular Plasma and Plunging Plasma

Abstract
This paper investigates the power and efficiency of energy extraction via magnetic reconnection in the Kerr-Sen-AdS4 black hole, focusing primarily on the circular orbit region and the plunging region. After introducing the black hole event horizon, boundary of the ergosphere, and photon sphere radius, we first study the energy extraction via magnetic reconnection in circular orbit region. We plot the allowed regions for energy extraction and present the energy extraction power and efficiency, comparing them with the Blandford-Znajek (BZ) mechanism. It was found that energy extraction is still possible even at a spin parameter of 0.5, which is lower than previous results, and the power can be higher than the BZ process in certain regions. In addition, we find the dilatonic scalar charge $b$ and the AdS radius $l$ contribute to lowering the spin threshold required for energy extraction. We also study the energy extraction in the plunging region by analyzing the allowed regions for energy extraction, and the power as well as efficiency. It was found that the energy extraction is possible even at a spin as low as 0.25. Especially, we find that parameter $b$ promotes lowering the energy extraction spin threshold while $l$ suppresses it, which is different from the circular orbit case. Moreover, the power and efficiency of energy extraction in the plunging region are higher than in the circular orbit region, indicating a stronger energy extraction capability in the plunging region.
I Introduction
Magnetic reconnection is crucial in astrophysics as it rapidly releases magnetic energy, generating significant power. It is often used to explain solar flares 1 , coronal mass ejections 2 and numerous other astronomical phenomena 3 ; 4 . Magnetic reconnection has been extensively studied in magnetohydrodynamics (MHD) 5 ; 6 ; 7 . Recent studies 8 ; 9 ; 10 indicate that spacetime curvature significantly influences the properties of magnetic reconnection, and the effect of gravity on the reconnection process requires further exploration. The analysis of polarized emission around the supermassive black hole in the center of M87* published by the Event Horizon Telescope collaboration 11 also confirms the presence of magnetic fields around black holes. This demonstrates the importance of analyzing magnetic reconnection processes within the framework of general relativity. Furthermore, magnetic reconnection serves as a mechanism for extracting energy from rotating black holes. Comisso and Asenjo first demonstrated in the Kerr black hole that the energy extraction power via this mechanism is typically higher than the traditional Blandford-Znajek mechanism 12 . Subsequently, this work was rapidly extended to other rotating black holes 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 ; 21 ; 22 ; 23 ; 24 ; 25 . However, this mechanism requires the black hole spin to be not too low. For a Kerr black hole, with fixed $\xi=\pi/12,\sigma=100$, the minimum spin is 0.85 12 . In Kerr-de Sitter spacetime, under the same conditions, the minimum spin is no less than 0.7 17 . In rotating regular black hole, under the same conditions and $k=0.1$, the minimum spin is about 0.79 18 . In rotating hairy black hole, under the same conditions and $\lambda=0.5,h_{0}=1.5$, the minimum spin is about 0.86 19 . All the results suggest that the required black hole spin for magnetic reconnection differs for different gravitational backgrounds. Exploring whether energy extraction via magnetic reconnection at lower spins in other gravitational spacetimes thus is a worthwhile topic. In this paper, we study magnetic reconnection in the Kerr-Sen-AdS4 black hole. We intend to investigate how the dilatonic scalar charge $b$ and AdS radius $l$ affect the magnetic reconnection process, and the power and efficiency. Especially whether this system is more effective, compared with previous results, to extract energy via magnetic reconnection. Our results show that the energy extraction is possible even at spins around 0.5 in this system, which is more effective than previous results. This reason stems from that both parameters $b$ and $l$ promote lowering the spin threshold for energy extraction.
Notably, the plasmas in the aforementioned studies all move on circular orbits outside the photon sphere radius. It is worth exploring whether the magnetic reconnection process can be studied in other regions, such as the plunging region. Especially, whether the power and efficiency of energy extraction is the same in different regions. Recently, Ref. 26 extended energy extraction to the plunging region. They found that the energy extraction efficiency in the plunging region is higher than in the circular orbit region, and the allowed region for energy extraction is also larger. In the Kerr-Sen-AdS4 black hole, we also study the magnetic reconnection process in the plunging region, analyzing the impact of the dilatonic scalar charge $b$ and AdS radius $l$ on energy extraction. We find that even at a spin of about 0.25, this black hole can still extract energy in the plunging region, which is lower than Kerr spacetime. We also find that, unlike circular orbits, in the plunging region, $b$ promotes lowering the energy extraction spin threshold, while $l$ suppresses it.
The remainder of this paper is organized as follows. Section 2 introduces the Kerr-Sen-AdS4 spacetime and its characteristics. Section 3.1 describes the magnetic reconnection process in circular orbits. Section 3.2 plots the parameter space for magnetic reconnection in circular orbits. Section 3.3 analyzes the power and efficiency of energy extraction in circular orbits. Section 4.1 introduces the magnetic reconnection process in the plunging region. Section 4.2 analyzes the efficiency and power in the plunging region. We present our conclusions in Section 5.
II Kerr-Sen-AdS4 Spacetime Introduction
In Boyer-Lindquist(BL) coordinates and geometric units $(c=G=1)$, the Kerr-Sen-AdS4 metric is given by 27 :
$ds^{2}$ $=\frac{-\Delta_{r}+a^{2}\Delta_{\theta}\sin^{2}\theta}{\rho^{2}}dt^{2}+\frac{\rho^{2}}{\Delta_{r}}dr^{2}+\frac{\rho^{2}}{\Delta_{\theta}}d\theta^{2}$ (1)
$+2\frac{a\sin^{2}\theta}{\rho^{2}\Xi}\left[\Delta_{r}-\Delta_{\theta}(r^{2}+2br+a^{2})\right]dt\,d\varphi$
$+\frac{\sin^{2}\theta}{\rho^{2}\Xi^{2}}\left[\Delta_{\theta}(r^{2}+2br+a^{2})^{2}-a^{2}\Delta_{r}\sin^{2}\theta\right]d\varphi^{2},$
where
$\Delta_{r}=\left(1+\frac{r^{2}+2br}{l^{2}}\right)\left(r^{2}+2br+a^{2}\right)-2Mr,$ (2)
$\Delta_{\theta}=1-\frac{a^{2}}{l^{2}}\cos^{2}\theta,\Xi=1-\frac{a^{2}}{l^{2}},\rho^{2}=r^{2}+2br+a^{2}\cos^{2}\theta,l^{2}=-\frac{3}{\Lambda},$ (3)
here $a=J/M$ is the angular momentum per unit mass, $b=Q^{2}/2M$ is the dilatonic scalar charge, $M$ is the black hole mass, $Q$ is the black hole charge, $\Lambda$ is the negative cosmological constant, $l$ is AdS curvature radius. When $b=0$, the metric reduces to the Kerr-AdS, when $l=\infty$, it reduces to the Kerr-Sen black hole, and when both conditions are satisfied, it reduces to the Kerr black hole.
The horizons of the Kerr-Sen-AdS4 black hole are determined by $\Delta_{r}=0$. This equation generally has two positive roots: the larger one is the event horizon, and the smaller one is the inner horizon. The boundary of ergosphere is determined by $g_{tt}=0$ , having one root outside the event horizon. We consider the equations of motion for particles on the equatorial plane ($\theta=\pi/2$) with $M=1$. The radial geodesic equation can be obtained from the Hamilton-Jacobi equation 28
$\left(\frac{dr}{d\lambda}\right)^{2}=-\Delta_{r}\left[K+\varepsilon\left(r^{2}+2br\right)\right]+\left[\left(r^{2}+2br+a^{2}\right)E-aL\Xi\right]^{2}=R(r),$ (4)
where $\lambda$ is Mino time 29 , related to proper time $\tau$ by $\frac{d\tau}{d\lambda}=\rho^{2},K$ is the Carter constant 30 , satisfying $K=(aE-L\Xi)^{2}$, $E$ is the energy, $L$ is the angular momentum. For photons, $\varepsilon=0$, for massive particles, $\varepsilon=1$. We consider particles on circular orbits in the equatorial plane. The radii of these circular orbits range from infinity down to the photon sphere. The photon sphere radius satisfies
$R(r)=0,R^{\prime}(r)=0.$ (5)
Typically, there are two solutions: one for prograde orbits (co-rotating with the black hole) and one for retrograde orbits (counter-rotating). In Fig. 1, we plot the photon sphere radius, boundary of ergosphere, and event horizon $r$ as functions of $a$. Green represents retrograde orbits, purple represents prograde orbits, blue represents the boundary of ergosphere, and red represents the event horizon. We can see that both the cosmological constant and dilatonic scalar charge reduce the maximum spin and the boundary of the ergosphere compared to the Kerr black hole. The maximum allowed spin for Kerr-Sen-AdS4 is less than 1, and the boundary of ergosphere is smaller than 2.
Figure 1: Photon sphere radius, boundary of ergosphere, and event horizon $r$ as functions of $a$. The top row shows changes in the four curves for fixed $l$ and varying $b$. The bottom row shows changes for fixed $b$ and varying $l$.
From the top row of Fig. 1, we see that for a fixed $l$, as $b$ increases, the allowed maximum spin decreases. This is crucial for enabling energy extraction at lower spins. From the bottom row, we see that for a fixed $b$, as $l$ decreases, the allowed maximum spin also decreases. Increasing $b$ and decreasing $l$ would increase the deviation from the Kerr black hole. The difference is that in the former case ($b$ increase), the positions of the boundary of ergosphere, event horizon, and photon sphere radius decrease, and the boundary of ergosphere doesn’t bend, while in the latter case ($l$ decrease), the positions of the boundary of ergosphere and event horizon decrease, and the boundary of ergosphere bends. Since energy extraction occurs between the event horizon and the boundary of ergosphere, we only need to consider prograde orbits suggested by Fig. 1. At the photon sphere radius, the particle’s energy and angular momentum approach infinity. Therefore, circular orbit energy extraction actually occurs between the photon sphere radius and the boundary of ergosphere. The Keplerian angular velocity for particles on prograde orbits is 28
$\Omega=\frac{-\partial_{r}g_{t\varphi}+\sqrt{\left(\partial_{r}g_{t\varphi}\right)^{2}-\left(\partial_{r}g_{tt}\right)\left(\partial_{r}g_{\varphi\varphi}\right)}}{\partial_{r}g_{\varphi\varphi}}.$ (6)
This formula comes from the definition of Keplerian angular velocity, $\Omega=\frac{d\varphi/d\tau}{dt/d\tau}$. It can be seen that this formula only involves the angular velocity in the azimuthal direction, not the radial direction, which is characteristic of circular orbits.
III Extracting Kerr-Sen-AdS4 Black Hole Energy in the Circular Orbit Region
III.1 Magnetic Reconnection Process in Circular Orbits
We use the Zero Angular Momentum Observer (ZAMO) frame for observation 31 . In the ZAMO frame, the metric takes the form
$ds^{2}=-d\hat{t^{2}}+\sum_{i=1}^{3}\left(d\hat{x^{i}}\right)^{2}=\eta_{\mu\nu}d\hat{x^{\mu}}d\hat{x^{\nu}},$ (7)
where
$d\hat{t}=\alpha dt,d\hat{x^{i}}=\sqrt{g_{ii}}dx^{i}-\alpha\beta^{i}dt,$ (8)
with
$\alpha=\left(-g_{tt}+\frac{g_{\varphi t}^{2}}{g_{\varphi\varphi}}\right)^{1/2},\beta^{\varphi}=\frac{\sqrt{g_{\varphi\varphi}}\omega^{\varphi}}{\alpha},\beta^{r}=\beta^{\theta}=0,\omega^{\varphi}=\frac{-g_{\varphi t}}{g_{\varphi\varphi}}.$ (9)
Therefore, in the ZAMO frame, the Keplerian velocity is
$\hat{v}_{K}=\frac{1}{\alpha}\left[\sqrt{g_{\varphi\varphi}}\Omega-\alpha\beta^{\varphi}\right].$ (10)
We adopt the single-fluid plasma approximation, with the energy-momentum tensor
$T^{\mu\nu}=pg^{\mu\nu}+wu^{\mu}u^{\nu}+F_{\sigma}^{\mu}F^{\nu\sigma}-(1/4)g^{\mu\nu}F^{\alpha\beta}F_{\alpha\beta},$ (11)
here $w,p,u,F$ are the plasma enthalpy density, pressure, four-velocity, and electromagnetic field tensor, respectively. Assuming the reconnection process is highly efficient, meaning magnetic energy is almost entirely converted to kinetic energy, we can neglect the electromagnetic field tensor part. Using the adiabatic and incompressible nature of the plasma, the energy at infinity per enthalpy for the accelerated and decelerated plasma is approximately 12
$e_{\pm}^{\infty}$ $=\frac{-\alpha g_{\mu 0}T^{\mu 0}}{w}$ (12)
$=\alpha\hat{\gamma}_{K}\left[\left(1+\beta^{\varphi}\hat{v}_{K}\right)\left(1+\sigma\right)^{1/2}\pm\cos\xi\left(\beta^{\varphi}+\hat{v}_{K}\right)\sigma^{1/2}-\frac{1}{4}\frac{\left(1+\sigma\right)^{1/2}\mp\cos{\xi}\hat{v}_{K}\sigma^{1/2}}{\hat{\gamma}_{K}^{2}\left(1+\sigma-\cos^{2}{\xi}\hat{v}_{K}^{2}\sigma\right)}\right],$
where $\xi$ is the azimuthal angle of the fluid in the local rest frame, $\sigma$ is the plasma magnetization parameter, and $\hat{\gamma}_{K}$ is the Lorentz factor of $\hat{v}_{K}$ , which can be expressed as
$\hat{\gamma}_{K}=(1-\left.\hat{v}_{K}^{2}\right)^{-1/2}.$ (13)
For a relativistically hot plasma, $w=4p$. Similar to the Penrose process 32 , energy extraction requires satisfying two conditions simultaneously, namely
$e_{-}^{\infty}<0,\Delta e_{+}^{\infty}=e_{+}^{\infty}-\left[1-\frac{\Gamma}{4(\Gamma-1)}\right]=e_{+}^{\infty}>0,$ (14)
here $\Gamma$ is the polytropic index, which we take as 4/3. We plot $e_{+}^{\infty}$ and $e_{-}^{\infty}$ in Figs. 2, 3, and 4. Here $r$ is the dominant reconnection point, called the X-point in 12 .
Figure 2: The variation of $e_{+}^{\infty}$ and $e_{-}^{\infty}$ with $\sigma$ under different azimuth angles, in which ${a}=0.98,{b}=0.01,l=20,{r}=1.2$.
From Fig. 2, it can be seen that $e_{+}^{\infty}$ is always greater than 0, while $e_{-}^{\infty}$ is almost always less than 0. Therefore, for energy extraction to occur, it mainly depends on whether $e_{-}^{\infty}$ is less than 0. As $\sigma$ increases, $e_{-}^{\infty}$ decreases. As the azimuthal angle increases, $e_{-}^{\infty}$ increases. This behavior is similar to the general case.
Figure 3: The variation of $e_{+}^{\infty}$ and $e_{-}^{\infty}$ with $\sigma$ under different $l$
From Fig. 3, we see that $e_{+}^{\infty}$ and $e_{-}^{\infty}$ do not vary monotonically with $l$, and their behavior depends on the location of the reconnection point. For smaller $r$, the difference for different $l$ is more obvious.
Figure 4: The variation of $e_{+}^{\infty}$ and $e_{-}^{\infty}$ with $\sigma$ under different $b$, in which ${a}=0.9,{r}=1.6,l=20,\xi=\pi/12$.
From Fig. 4, we see that as $b$ increases, $e_{-}^{\infty}$ gradually decreases, indicating that energy extraction becomes more favorable. It should be note that $a$ cannot be set to 0.98 here, because $b=0.04$ and $b=0.08$, a naked singularity appears.
III.2 Parameter Space for Energy Extraction Via Magnetic Reconnection in Circular Orbits
We plot the allowed region for energy extraction ($e_{-}^{\infty}<0$) in the $r-a$ plane in Figs. 5-8. The red solid line represents the event horizon, the blue solid line represents the ergosphere, the purple dashed line represents the photon sphere radius, and from left to right, $\sigma=100$, 30, 10, 3.
Figure 5: Allowed energy extraction regions for different $b$.
From Fig. 5, we see that as $\sigma$ increases, the allowed energy extraction region also increases, which is consistent with 12 . As $b$ increases, the maximum allowed spin decreases, enabling energy extraction at lower spins. For example, the maximum spin allowed for the black hole in Fig. 5(a) is 0.98753, while in Fig. 5(c) it is 0.6977. The minimum spin allowing energy extraction also decreases, from 0.85 in Fig. 5(a) to 0.64 in Fig. 5(c). Simultaneously, due to the decrease in the ergosphere, event horizon, photon sphere radius, and the position of the reconnection layer also decreases.
Figure 6: Allowed energy extraction regions for different $l$.
From Fig. 6, we see that as $l$ decreases, the maximum allowed spin also decreases, enabling energy extraction at lower spins. The minimum spin allows energy extraction decreases too, analogous to the effect of increasing $b$. Therefore, we conclude that in circular orbits, both parameters $b$ and $l$ promote lowering the spin threshold for energy extraction.
Figure 7: Allowed energy extraction regions for different $\xi$.
From Fig. 7, we see that as $\xi$ decreases, the area of the allowed energy extraction region increases, the minimum allowed spin decreases, and the position of the reconnection layer increases. Thereafter, we will not pay attention to the effect of $\xi$ on the magnetic reconnection. Without loss of generality, we choose $\xi=\pi/12$.
Figure 8: Allowed energy extraction region for ${b}=0.36,l=2,\xi=\pi/12$
In Fig. 8, we intend to explore how low for the spin $a$ can be as the magnetic reconnection occurs. We find for the large $b$ and small $l$, the spin will be lower. The result shows that the allowed spin range for energy extraction is between 0.51 and 0.521, demonstrating that energy extraction is possible at moderate spins.
III.3 Power and Efficiency of Energy Extraction in Circular Orbits
Having established the feasibility of energy extraction, we compare the power and efficiency. The energy extraction power is 12
$P=-e^{\infty}_{-}wA_{in}U_{in}.$ (15)
[…]
$A_{in}\sim\left(r_{E}^{2}-r_{ph}^{2}\right),$ (16)
where $r_{E}$ is the boundary of ergosphere and $r_{p{h}}$ is the photon sphere radius.
Figure 9: Energy extraction power for different $\sigma$, in which ${b}=0.01,l=20,\xi=\pi/12,{a}=0.98$.
We plot the energy extraction power per enthalpy density $P/w$ with respect to $r$ in Figs. 9-12. Red dashed, purple dashed, and green dashed lines represent the event horizon, photon sphere, and boundary of ergosphere, respectively. Red solid, orange solid, blue solid, and green solid curves represent $\sigma=3,~{}10,~{}30,~{}100$.
Figure 10: Energy extraction power for different $\sigma$, in which ${b}=0.3,l=20,\xi=\pi/12,{a}=0.696$.
From Fig. 9 we can see that the power starts outside the photon sphere radius, consistent with circular orbit characteristics. The power first increases and then decreases. As $\sigma$ increases, the power increases. Comparing Figs. 9 and 10, we see that as $b$ increases, the energy extraction power decreases for the same $\sigma$. Note that in Figs. 9 and 10, $a$ cannot be the same because the allowed spin of the black hole varies under different $b$ values. This is similar to 19 , in which $a$ cannot be the same because the allowed spin of the black hole varies under different $h_{0}$ values. Comparing Figs. 9 and 11, we see that as $l$ decreases, the energy extraction power also decreases due to the lower spin. Similarly, $a$ cannot be the same because the allowed spin of the black hole varies under different $l$. Especially, for some special values of $b$ and $l$, which is shown in Fig. 12, we find the energy extraction power becomes even lower.
Figure 11: Energy extraction power for different $\sigma$, in which ${b}=0.01,l=2,\xi=\pi/12,{a}=0.856$.
Figure 12: Energy extraction power for different $\sigma$, in which ${b}=0.36,l=2,\xi=\pi/12,{a}=0.52$. $b$ and $l$.
Next, we intend to discuss the energy extraction efficiency, which is defined as 12
$\eta=\frac{e_{+}^{\infty}}{e_{+}^{\infty}+e_{-}^{\infty}}.$ (17)
Since the energy extraction condition requires $e_{-}^{\infty}<0,e_{+}^{\infty}>0$, for energy extraction to occur, the efficiency must be greater than 1. We plot the energy extraction efficiency with respect to $r$ in Figs. 13-16. We fix $\xi=\pi/12,\sigma=100$.
Figure 13: Energy extraction efficiency for different $a$, in which ${b}=0.01,l=20$.
Figure 14: Energy extraction efficiency plot for different $a$, in which ${b}=0.3,l=20$.
In Fig. 13, we plot the efficiency for different $a$ under the small $b$. Like the power, we find the efficiency first increases and then decreases. As $a$ increases, the efficiency increases. For the larger $b$, the efficiency for different $a$ is shown in in Fig. 14. Comparing Figs. 13 and 14, we see that as $b$ increases, the energy extraction efficiency decreases due to the lower spin, similar to the power trend. In Figs. 13 and 14, $a$ cannot be the same because the allowed spin of the black hole varies under different $b$ values. As $b$ is fixed, we also can discuss the effect of $l$ on the energy extraction efficiency. Comparing Figs. 13 and 15, we see that as $l$ decreases, the energy extraction efficiency also decreases due to the lower spin, similar to the power trend. Especially, for some special values of $b$ and $l$, which is shown in Fig. 16. Comparing From Fig. 13 and From Fig. 16, we find the efficiency becomes lower, which is also similar to the power trend.
Figure 15: Energy extraction efficiency plot for different $a$, in which ${b}=0.01,l=2$.
Figure 16: Energy extraction efficiency for different $a$, in which ${b}=0.36,l=2$.
Finally, we compare the energy extraction power ratio. We compare it with the very popular Blandford-Znajek (BZ) mechanism 36 . The BZ energy extraction power is 37
$P_{BZ}=\frac{\kappa}{16\pi}\Phi_{H}^{2}\left(\Omega_{H}^{2}+c_{1}\Omega_{H}^{4}+c_{2}\Omega_{H}^{6}+\mathcal{O}\left(\Omega_{H}^{8}\right)\right),$ (18)
where $\kappa=0.05,c_{1}=1.38,c_{2}=-9.2$ are numerical constants, and $\Phi_{H}$ is the magnetic flux threading one hemisphere of the black hole event horizon, which takes the form
[…]
where $r_{+}$ is the event horizon radius, $B_{0}=\sqrt{w\sigma},~{}\Omega_{H}$ is the angular velocity of the event horizon, namely
$\Omega_{H}=\left.\frac{-g_{\varphi t}}{g_{\varphi\varphi}}\right|_{r=r_{+}}=\frac{a\Xi}{r_{+}^{2}+2br_{+}+a^{2}}.$ (20)
Therefore, the power ratio is
$\frac{P}{P_{BZ}}=\frac{-4e^{\infty}_{-}A_{in}U_{in}\Xi^{2}}{\kappa\pi\sigma\left(\Omega_{H}^{2}+c_{1}\Omega_{H}^{4}+c_{2}\Omega_{H}^{6}\right)\sin^{2}(\xi)\left(r_{+}^{2}+2br_{+}+a^{2}\right)^{2}}.$ (21)
From (21), it can be seen that as long as the azimuthal angle is sufficiently small, the power ratio can always be made greater than 1, meaning the magnetic reconnection mechanism always has higher power than the BZ mechanism. Without loss of generality, we fix $\xi=\pi/12$. Taking parameters for high spin (${b}=0.01,l=20,\sigma=3,{a}=0.98,{r}=1.2$), we obtain a power ratio of 10.3056 > 1, confirming that magnetic reconnection power exceeds BZ power. For example, at low spin and large deviation (${b}=0.36,l=2,{a}=0.52,{r}=0.5,\sigma=7$), the power ratio is 5.56032. Even though the absolute power and efficiency are reduced in this case, the magnetic reconnection power still exceeds the BZ power. This is thanks to the factors $b$ and $l$, which lower the spin threshold and reflect the high efficiency of magnetic reconnection.
IV Extracting Kerr-Sen-AdS4 Black Hole Energy in the Plunging Region
IV.1 Magnetic Reconnection Process in the Plunging Region
Previous studies considered plasma moving on circular orbits outside the photon sphere radius. Here we consider another scenario as in 26 . That is, the plasma initially moves on a circular orbit outside the Innermost Stable Circular Orbit (ISCO), then plunges inward starting from the ISCO. Circular orbits inside the ISCO are unstable, any perturbation induces radial velocity, initiating the plunge. The region $r<r_{I}$ (where $r_{I}$ is the ISCO radius, typically larger than the photon sphere radius) is called the plunging region 38 . In the plunging region, because there is radial velocity, the Keplerian velocity formula in (12), which only contained the azimuthal component, is no longer applicable. We still use the convenient ZAMO frame. The relationship between the four-velocity in the ZAMO frame and the BL frame is
$\hat{U}^{\mu}=\hat{\gamma}_{s}\left\{1,\hat{v}_{s}^{(r)},0,\hat{v}_{s}^{(\varphi)}\right\}=\left\{\frac{E-\omega^{\varphi}L}{\alpha},\sqrt{g_{rr}}U^{r},0,\frac{L}{\sqrt{g_{\varphi\varphi}}}\right\},$ (22)
where
$\left(U^{r}\right)^{2}=\left(\frac{dr}{d\tau}\right)^{2}=\frac{R(r)}{\rho^{4}},$ (23)
$\rho^{2}\left(\frac{dt}{d\tau}\right)=\frac{E\left(r^{2}+2br+a^{2}\right)^{2}-a\left(r^{2}+2br+a^{2}\right)L\Xi}{\Delta_{r}}-\frac{\sin^{2}\theta}{\Delta_{\theta}}\left(a^{2}E-\frac{aL\Xi}{\sin^{2}\theta}\right),$ (24)
$E=-\left(g_{tt}+\Omega g_{t\varphi}\right)\left(\frac{dt}{d\tau}\right),$ (25)
$L=\left(g_{t\varphi}+\Omega g_{\phi\phi}\right)\left(\frac{dt}{d\tau}\right),$ (26)
The conserved quantities in the plunging region are the energy $E$ and angular momentum $L$, at the ISCO, we label them as
$E_{I}=E\left(r_{I}\right),L_{I}=L\left(r_{I}\right).$ (27)
Substituting (27) into (23), and using $R(r)$ satisfying (4) with $\varepsilon=1$, we get
$U^{r}=-\frac{1}{\rho^{2}}\sqrt{-\Delta_{r}\left[\left(aE_{I}-L_{I}\Xi\right)^{2}+\left(r^{2}+2br\right)\right]+\left[\left(r^{2}+2br+a^{2}\right)E_{I}-aL_{I}\Xi\right]^{2}}.$ (28)
The negative sign indicates inward motion. Substituting (28) into (22) gives $\hat{v}_{s}^{(r)},\hat{v}_{s}^{(\varphi)}$, and
$\hat{v}_{s}=\sqrt{\left(\hat{v}_{s}^{(r)}\right)^{2}+\left(\hat{v}_{s}^{(\varphi)}\right)^{2}}$,$\hat{\gamma}_{s}$ is the Lorentz factor of $\hat{v}_{s}$ , $r_{I}$ satisfies (5) and also $R^{\prime\prime}(r)=0$. In this case, the $e_{\pm}^{\infty}$ of (12) becomes as 39
$e_{\pm}^{\infty}$ $=\alpha\hat{\gamma}_{s}\gamma_{\text{out }}\left[\left(1+\beta^{\varphi}\hat{v}_{s}^{(\varphi)}\right)\pm v_{\text{out }}\left(\hat{v}_{s}+\beta^{\varphi}\frac{\hat{v}_{s}^{(\varphi)}}{\hat{v}_{s}}\right)\cos\xi\mp v_{\text{out }}\beta^{\varphi}\frac{\hat{v}_{s}^{(r)}}{\hat{\gamma}_{s}\hat{v}_{s}}\sin\xi\right]$ (29)
$-\frac{\alpha}{4\hat{\gamma}_{s}\gamma_{\text{out }}\left(1\pm\hat{v}_{s}v_{\text{out }}\cos\xi\right)},$
where $v_{\text{out }}$ is the outflow speed and $\gamma_{\text{out }}$ is its Lorentz factor, which can be expressed as
$v_{\text{out }}=\sqrt{\frac{\sigma}{\sigma+1}},\gamma_{\text{out }}=\sqrt{1+\sigma}.$ (30)
Following the approach for circular orbits, we plot the allowed energy extraction region in the ${r}-{a}$ plane for the plunging case in Figs. 17-20. The red line represents the event horizon, blue line represents the ergosphere, purple line represents the photon sphere radius, and black line represents the ISCO. From left to right, $\sigma=100,30,10,3$.
Figure 17: Allowed energy extraction regions for different $b$ in the plunging region.
From Fig. 17(a), we see that as $\sigma$ increases, the allowed energy extraction region also increases. Comparing with Fig. 5(a), for $r>r_{I}$ , the allowed region matches the circular orbit case. For $r<r_{I}$, the allowed region in the plunging region is much larger than in circular orbits. Additionally, the minimum spin allowing energy extraction in the plunging region is lower than in circular orbits, and the position of the reconnection point is higher. This is consistent with 26 . For example, taking $\sigma=100$, the minimum spin for energy extraction in circular orbits (Fig. 5(a)) is 0.85, while in the plunging region, it is 0.31. Comparing the three panels of Fig. 17, we see that as $b$ increases, the minimum spin allowing energy extraction also decreases: from 0.31 in Fig. 17(a) to 0.25 in Fig. 17(c). This enables energy extraction from slowly rotating black holes.
Figure 18: Allowed energy extraction regions for different $l$ in the plunging region.
From Fig. 18, we see that as $l$ decreases, the minimum spin allowing energy extraction increases. This is inconsistent with the trend in the circular orbit region. Therefore, we conclude that in the plunging region, $b$ promotes lowering the energy extraction spin threshold, while $l$ suppresses it. However, even with this, the minimum spin in the plunging region is still lower than in circular orbits. For example, taking $\sigma=100,l=2$ (Fig. 18(c)), the minimum plunging region spin is 0.54, while for circular orbits (Fig. 6(c)), it is 0.795. This highlights the advantage of the plunging region.
Figure 19: Allowed energy extraction regions for different $\xi$ values in the plunging region.
From Fig. 19, we see that the shape of the allowed region changes with the azimuthal angle, which is a characteristic feature of the plunging region azimuthal dependence, which is consistent with 26 .
Figure 20: Allowed energy extraction region for ${b}=0.36,l=2,\xi=\pi/12$ in the plunging region.
Following Fig. 8, we plot the allowed region for large $b$ and small $l$ in Fig. 20. Here, the minimum spin for energy extraction at $\sigma=100$ is 0.375. If $l$=20, then $b=0.36>0.3$ would leads to a spin less than 0.25 based on previous patterns, but this is not the case here because the presence of $l$ raises the spin threshold.
IV.2 Power and Efficiency of Energy Extraction in the Plunging Region
Similarly, we plot the energy extraction power in Fig. 21, still using (15). To highlight the difference, we plot power for both plunging and circular orbit cases within $r<r_{I}$. We take ${b}=0.01,l=20,\xi=\pi/12,{a}=0.9,\sigma=100$, here all $r$ values are less than ISCO.
Figure 21: Red dashed, purple dashed, green dashed lines represent event horizon, photon sphere radius, ergosphere, respectively. Red solid line represents the circular orbit power and blue solid line represents the plunging region power.
From Fig. 21, it can be seen that the energy extraction power in the plunging region is higher than that in the circular orbit region. Similarly, we compare the energy extraction efficiency for both cases using (17), shown in Fig. 22. We use the same parameters in Fig. 21 and Fig. 22, that is ${b}=0.01,l=20,\xi=\pi/12,{a}=0.9,\sigma=100$.
Figure 22: Red dashed, purple dashed, green dashed lines represent event horizon, photon sphere radius, ergosphere, respectively. Red solid line represents the circular orbit efficiency and blue solid line represents the plunging region efficiency.
From Fig. 22, the energy extraction efficiency in the plunging region is higher than in the circular orbit region. Comparing with Fig. 21, we see that for the same parameters, the trend of power change is similar to the trend of efficiency change.
Finally, we calculate the power ratio for the plunging region compared to the BZ process, still using (21). Using the same parameters as the circular orbit high spin case (${b}=0.01,1=20,\sigma=3,{a}=0.98,{r}=1.2$, $\xi=\pi/12$), and noting that $r_{I}=1.379$ > $r$ satisfies the plunging region condition, we get a power ratio of 15.6634 > 10.3056. This again shows that plunging region power exceeds both circular orbit power and BZ power. Now considering a low-spin case based on Fig. 17(c) (${b}=0.3,1=20,\xi=\pi/12$), spin $a$ can be 0.3 or lower. However, note that here the photon sphere radius is above the ergosphere, so (16) is no longer valid and should be replaced by
$A_{in}\sim\left(r_{E}^{2}-r_{+}^{2}\right).$ (31)
Taking $\sigma=100,{a}=0.3,{r}=1.34,\xi=\pi/12$, we get a power ratio of 0.0393679, which is very low. This is understandable because in the plunging region, to achieve low spins, $\sigma$ must be large. According to (21), $\sigma$ appears in the denominator, reducing the power ratio. Nevertheless, the ability to extract energy via magnetic reconnection at spins as low as 0.25 is very interesting, for previous studies rarely achieved such low spins.
V Conclusion
Magnetic reconnection provides a new mechanism for extracting energy from rotating black holes. How to extract energy efficiently and conveniently is the goal pursued by physicists. Centering on this goal, we conducted an investigation in Kerr-Sen-AdS4 black hole. We discussed not only in the circular orbit region but also in the plunging region. First, we introduced the event horizon, boundary of ergosphere, and photon sphere radius of this black hole. We found that due to the presence of the dilatonic scalar charge $b$ and AdS radius $l$, the maximum allowed spin is reduced compared to the Kerr black hole. Then, we described the energy extraction mechanism via magnetic reconnection in circular orbits. We analyzed the feasibility of energy extraction for this black hole, plotted the parameter space, and analyzed the influence of parameters $b$, $l$, $\xi$ and $\sigma$ on energy extraction. We also analyzed the energy extraction power and efficiency, especially under the influence of $b$ and $l$, and compared them with the BZ mechanism, finding that magnetic reconnection power can exceed BZ power. Finally, we studied energy extraction in the plunging region, describing the magnetic reconnection mechanism there. We continued analyzing the influence of parameters $b$, $l$, $\xi$ and $\sigma$, comparing the results with the circular orbit case. We found that the power and efficiency of energy extraction are higher in the plunging region.
For circular orbits, we found that energy extraction is possible even at a spin of 0.5. For the plunging region, we found that energy extraction is possible even at a spin of 0.25. This is caused by the influence of the two parameters $b$ and $l$ in this black hole. We also found that in circular orbits, both $b$ and $l$ promote lowering the spin threshold for energy extraction. However, in the plunging region, the situation is different, that is, $b$ promotes lowering the spin threshold, while $l$ suppresses it.
This conclusion indicates that the plunging region has a stronger energy extraction capability. To extract energy, it is best to operate in the plunging region. For circular orbits, increasing $b$ or decreasing $l$ enables energy extraction at lower spins. For the plunging region, increasing $b$ or increasing $l$ enables energy extraction at lower spins.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grants No. 12375043).